OpenStudy (anonymous):

Simplify. Assume no radicands were formed by raising negative quantities to even powers.

OpenStudy (anonymous):



OpenStudy (anonymous):

What is square root of 9??

OpenStudy (anonymous):

3

OpenStudy (anonymous):

\(\huge \color{green}{\textbf{Welcome To Openstudy..}}\)

OpenStudy (anonymous):

Yes you are right.. And what is the square root of 16??

OpenStudy (anonymous):

4. What do you do with the exponents?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

Oh alright, what do you do if they aren't the same?

OpenStudy (anonymous):

Firstly we solve for numerator.. \(x^7 = x^6 \times x\) Yes or no??

OpenStudy (anonymous):

Wait, where did the \[\chi ^{6}\] come from?

OpenStudy (anonymous):

See, How can you write x^3?? \(x^3 = x \times x \times x\) Yes or no??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, how can you write \(x^7\) ??

OpenStudy (anonymous):

\[x \times x \times x \times x \times x \times x \times x\]

OpenStudy (anonymous):

Can I write first 6 x as \(x^6\) ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So what you are left with?? \(x\) So, \[x^7 = x \times x \times x \times x \times x \times x \times x\] \[x^7 = x^6 \times x\] Getting?

OpenStudy (anonymous):

yeah I got it now :)

OpenStudy (anonymous):

What is the square root of \(x^6\) ??

OpenStudy (anonymous):

Can you solve that? 6 can't be reduced? I'm not really sure actually.

OpenStudy (anonymous):

See, square root of \(x^2\) is \(x\)... Square root of \(x^4 = x^2\) So what will be the square root of x^6 ??

OpenStudy (anonymous):

Would it be \[x^3\] then?

OpenStudy (anonymous):